A fathom is a unit of length, usually reserved for measuring the depth of water. A fathom is approximately 6 ft in length. Take the distance from Earth to the Moon to be 252,000 miles, and use the given approximation to find the distance in fathoms. (Enter a number with the correct number of significant figures.)
Answer #1

We know that, 1 mile = 5280 ft

Converting the distance from earth to moon to ft,

=> d = 252000 × 5280 ft = 1,330,560,000 ft

Given that, 1 fathom = 6 ft

Therefore the distance in fathoms will be,

=> d = 1,330,560,000/6 = 221,760,000 fathoms

Therefore the distance between earth and moon is 221,760,000 fathoms.